40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog, Page 57 of 58  Not logged in ELOG logo
ID Date Author Type Category Subject
  69   Sat Jan 29 00:27:42 2011 DmassMiscPlotsSilicon - absorption plots

Is the X-axis of the third figure wavelength in nm?

  68   Fri Jan 28 21:05:01 2011 FrankMiscPlotsSilicon - absorption plots

reason: Does it make sense to use silicon substrates for 1064nm coated cavity mirrors?

some plots i've found (nothing for cryo temps so far):

abs_Si_1064nm_p_dopant.png

 

abs_Si_1064nm_n_dopant.png

 

Absorption_edge_of_silicon.png

 

absorption_silicon.png

  67   Fri Jan 28 20:23:37 2011 DmassLaserSchematicsDatasheets

Here are the datasheets for both lasers and the ATF mirrors

Attachment 1: ATFCoatingsResized.png
ATFCoatingsResized.png
Attachment 2: EmcoreDatasheetResized.png
EmcoreDatasheetResized.png
Attachment 3: CovegaDatasheets.png
CovegaDatasheets.png
  66   Fri Jan 28 17:53:25 2011 FrankLab InfrastructureGenerallab cleaned

cleaned the lab today. Particle count went up by more than a factor of 10 while cleaning. Will post latest data later today.
But this indicates that there is still a lot of dust in the lab which we stir up when actually working in the lab.
So we should clean more often in the near future to get rid of the dust before we install all our optics on the table.
Current cleaning of the floor and the cupboards/racks/desks does not take more than 10min with one person.
As we have a super clean air supply and good overpressure this should bring us down very quickly to a real clean lab without extensive cleaning.
The lamps hanging from the ceiling are very dirty so we should schedule a day where we clean those once. There is a thick dust layer on them.

  65   Thu Jan 27 17:50:36 2011 FrankHowToCryocheap cooler down to 77K

While browsing the web for cryo-stuff i've found something interesting:
A company called Superconducting Technology is producing low noise amplifiers and filters for wireless communication.
The way they build the "low-noise" part is by cooling it down to 77K using a Stirling Cryocooler. It contains a cryocooler that is rated at 140 watts of input power, and is extensively documented here (http://books.google.com/books?id=POLgG5mma6IC&pg=PA75). Those devices (the entire HF filter unit) are sold for ~$300 on Ebay:

Superfilter

more information here:

http://www.neiu.edu/~pjdolan/Link4/

http://benkrasnow.blogspot.com/2008/08/diy-liquid-nitrogen-generator.html

http://www.electronics-related.com/usenet/design/show/232736-1.php

http://www.flickr.com/photos/kc6qhp/sets/72157625382973371/

 

If i find one in the future i will buy one :-)

  64   Thu Jan 27 14:11:50 2011 FrankMiscPlotsReference Cavity HOM spacing

The ATF stock values are for fused silica. We should ask ATF about standard radii of curvature for silicon substrates, if there are any.

Quote:

Here are some plots of higher order mode spacing using the following parameters:

  • 4 inch cavity
  • ATF stock curvatures (0.5m, 1m, 2m)
  • Symmetric cavity
  • Finesse = 10,000

I plotted the first 10, but have only shown those which are close to the 00 here.

For R1=R2=0.5m (red x's, which seem fine):

  • Rayleigh range = 15 cm
  • waist = 272 um
  • Spot size at mirrors = 273 um
  • cavity g-factor = 0.8

 

One idea which has been discussed is making a flat / curved cavity with hopes of measuring coating thermal noise by having a waist on one of the cavity mirrors. To reap real benefits from this we would want a Rayleigh range at least comparable to cavity length, preferably larger...That means using a mirror faster than 0.5m (which according to Frank is faster than ATF stock).

 

  63   Thu Jan 27 03:38:14 2011 dmassMiscPlotsReference Cavity HOM spacing

Here are some plots of higher order mode spacing using the following parameters:

  • 4 inch cavity
  • ATF stock curvatures (0.5m, 1m, 2m)
  • Symmetric cavity
  • Finesse = 10,000

I plotted the first 10, but have only shown those which are close to the 00 here.

For R1=R2=0.5m (red x's, which seem fine):

  • Rayleigh range = 15 cm
  • waist = 272 um
  • Spot size at mirrors = 273 um
  • cavity g-factor = 0.8

 

One idea which has been discussed is making a flat / curved cavity with hopes of measuring coating thermal noise by having a waist on one of the cavity mirrors. To reap real benefits from this we would want a Rayleigh range at least comparable to cavity length, preferably larger...That means using a mirror faster than 0.5m (which according to Frank is faster than ATF stock).

Attachment 1: RefCav1HOM.pdf
RefCav1HOM.pdf
  62   Tue Jan 25 11:56:20 2011 FrankUpdateDrawingsinitial cavity test setup

1550nm-setup_initial_setup_v2.png

  61   Mon Jan 24 23:17:30 2011 FrankUpdateCavitycavity locked (for a short time)

can lock it  for a second or so before loosing lock, but didn't optimize loop parameters.Still having problems getting sidebands so this is currently No 1 on my priority list.
I actually don't know what i'm locking to, but can lock to individual modes. It looks like there is kind of a zero crossing, but not really.

Using Zach's resonant EOM at 33MHz at the moment. Had to switch PD from Thorlabs PDA10CS to the 2GHz diode.
With a 13dBm mixer and 23dBm into the resonant EOM the "signal" is only about 7mV large and does not like a n error signal at all.
More like some funny oscillation. Probably due to the high frequency noise. Also tried other frequencies, mixers, photodiodes. Needs more investigation.

Modematching into the cavity is only about 25%. Could be either a non impedance matched cavity or our calculation for the mode matching is wrong. Will check this tomorrow.
Nevertheless this can't explain the non-existing error signals.

  60   Mon Jan 24 16:40:35 2011 FrankUpdateHVACparticle count - updated graph for the last week

CryoLab_data_plot_detail2.png

  59   Mon Jan 24 01:09:58 2011 FrankLab InfrastructureHVAClatest particle counts

History: When i've first measured the air quality in the cryo lab the particle count was several million. So i've installed a HEPA filter and the counts dropped down to several tens of thousands.
But the counts really didn't get much smaller even after mopping the lab several times. So i kept monitoring the lab over xmas to see what happens when noone is working in the lab for days.

When i came back this year the counts were still very high, even over xmas holidays. So i started searching. The filtered air had still a count of zero, so where did/does the dirt come from?
I realized that the lab doesn't have much overpressure compared to the other labs, probably due to the installation of the filter which creates a large pressure drop.
So i carefully checked the lab and finally simply blocked the sucking hole. Now we have plenty of overpressure in the lab and the counts went down.

Here the latest plot. I've fixed the sucking problem during the time without data a couple of days ago.

CryoLab_data_plot2.png

i will keep an eye on this and will reopen the sucking hole at least a little bit to see how much we have to block to maintain enough over pressure to keep the lab clean.

  58   Mon Jan 24 00:36:29 2011 FrankComputingGenerallab computer and network structure.

i've set up an old computer in the lab with Centos installed. User names and passwords are as usual.
Right now the computer is in the Caltech network only as we don't have a LIGO network connection in that lab.
The computer has a floppy drive installed to download data from scopes/analyzers.

The current public IP-address is 131.215.123.221 (dynamic, so might change in the near future)

Once we updated the wireless router in the ATF i would like make some minor changes to the network structure.
Suggested changes are changing the network mask from 255.255.255.0 to 255.255.0.0 on current computers in the ATF and PSL lab.
This would give us an easier network structure with the same/current subnetworks (10.0.0.x, 10.0.1.x, etc.), having access from any lab to any other.
Labs like TCS would get their own subnetwork 10.0.2.x and the Cryo/SUS lab 10.0.3.x.
Currently only the TCS computers (four or so) need to change the ip-addresses too.
The network mask can be changed whenever maintenance is due. Everything would still work even with the old netmask.
Non updated computers would have access limitations as they have right now until changed. 

I would like to install one of the wireless bridges (TEW-430APB) in the Cryo lab to get a link to the ATF. I've already tested that and it is working fine. So we don;t have to deal with another router/gateway.

  57   Sun Jan 23 23:24:01 2011 FrankUpdateCavityfirst setup for cavity test

below a picture of the first setup using one fiber coupled laser diode (right) mode matched to the FS cavity (left). First modes can be seen in transmission of the cavity. Further optimization required.

1550nm-setup.jpg

  56   Sun Jan 23 23:18:51 2011 FrankPhotosCavity1550nm cavity assembled

Peter and i assembled the first 1550nm cavity the day before i left for vacation. We used the spare LIGO reference cavity FS spacer and the ULE mirror set from ATF.
The mirrors are optically contacted to the spacer. Below some pictures...

100_1183.JPG

fused silica spacer

 

ule-mirrors.jpg

ULE mirror set: TOP:flat mirror  BOTTOM:curved mirror (R=0.5m)

 

100_1191.JPG

using first contact to clean spacer surface

 

100_1197.JPG

mirror with some dust on it

 

P1670661.JPG

cavity with mirror contacted

 

P1670702.JPG

assembled cavity

 

P1670686.JPG

assembled cavity

  55   Fri Jan 7 02:07:00 2011 DmassLaserLaserDriver Noise Measured

I kept the same sensitivity for both measurements, and checked with a 50 Ohm terminator that the noise floor was below the measurement. For the higher frequencies, the SNR may have gotten as small as 2 at high frequencies, but the measurement was never noise limited.

The spectra were noticably nonstationary in some not-obviously-repeating way, and I assume this has more to do with the lack of lining up. There was a lot of power at lower frequencies in several of the measurements, I retook a few.

  54   Thu Jan 6 23:32:26 2011 ZachLaserLaserBeat Found

I think your higher-frequency ITC 510 data is just the same low-frequency data vector plotted against the higher frequencies. Might want to double check that.

DYM: Zach Wins.

Quote:

Here is the RIN of the Diodes, as measured on the HP35670A.

A second plot is included, which shows what we think is the Driver contribution to the RIN...

  1. I took each laser diodes P vs I curve to get dP/dI, then divided by P_laser to get dRin/dI
  2. This gave me 0.138 W/A and 0.26 W/A for the Emcore and Covega Laser Didoes, respectively
  3. This corresponds to a dRIN/dI of 3.5/Amp and 5.2/Amp (if what I described in 1 is OK)
  4. The contributions to RIN in the noisebudget from the driver noise are HIGHER than the actual RIN measured
  5. I think this is fine and it merely means that the transfer function of current noise into diode RIN is not the flat unity gain model implicitly assumed in 1 above
  6. Maybe a good measurement to do at some point is to T a (low noise) buffer onto the input current to the diode, and measure the coherence and transfer function between the driver current noise and the laser diode RIN
  7. Using each laser diode's P(I) curve to get the transfer function of current noise into RIN seems increasingly less bad at low frequencies, which is I would expect (no crazy low frequency features in the transfer function) - and it looks like we might be limited by the driver noise for low frequency RIN.
  8. If RIN and frequency noise are coherent, and we close a loop on frequency noise, this won't be a huge problem, since actuator noise is supressed by our (not yet existent) loop gain, and it should crush the current noise...the question, as ever, is "Can we obtain enough bandwidth to lower the PSD to what we define as an acceptable value"?

 

Happy New Year CryoLab.

 

 

  53   Thu Jan 6 13:53:15 2011 FrankLaserLaserBeat Found

Hmm,you should check your measurements. If you measure the noise floor of your equipment and it is higher than your measurements taken obviously something must be wrong.

Also, the projected RIN from the current noise should match the measured RIN (or be lower). The TF for the laser diodes is almost flat and not -20dB at 100Hz or so. I don't see any physical reason why it should be so at low frequencies. To verify that simply use the current mod input of one of the drivers (check max mod range, might be only 10kHz or so) and measure the TF up to 10 or 100Hz and you will see that it is flat. If you want to be sure measure the TF using your shunts first.

Quote:

Here is the RIN of the Diodes, as measured on the HP35670A.

A second plot is included, which shows what we think is the Driver contribution to the RIN...

  1. I took each laser diodes P vs I curve to get dP/dI, then divided by P_laser to get dRin/dI
  2. This gave me 0.138 W/A and 0.26 W/A for the Emcore and Covega Laser Didoes, respectively
  3. This corresponds to a dRIN/dI of 3.5/Amp and 5.2/Amp (if what I described in 1 is OK)
  4. The contributions to RIN in the noisebudget from the driver noise are HIGHER than the actual RIN measured
  5. I think this is fine and it merely means that the transfer function of current noise into diode RIN is not the flat unity gain model implicitly assumed in 1 above
  6. Maybe a good measurement to do at some point is to T a (low noise) buffer onto the input current to the diode, and measure the coherence and transfer function between the driver current noise and the laser diode RIN
  7. Using each laser diode's P(I) curve to get the transfer function of current noise into RIN seems increasingly less bad at low frequencies, which is I would expect (no crazy low frequency features in the transfer function) - and it looks like we might be limited by the driver noise for low frequency RIN.
  8. If RIN and frequency noise are coherent, and we close a loop on frequency noise, this won't be a huge problem, since actuator noise is supressed by our (not yet existent) loop gain, and it should crush the current noise...the question, as ever, is "Can we obtain enough bandwidth to lower the PSD to what we define as an acceptable value"?

 

Happy New Year CryoLab.

 

 

  52   Thu Jan 6 09:56:00 2011 FrankLaserLaserDriver Noise Measured

Nice, but did you change the input sensitivity during the measurements or does the spectrum fluctuate a lot?
The noise spectra taken for different frequencies should line up, but especially for the ITC510 they don't.
You should check that again with some more overlap to see if you really measure the driver noise or only the preamp/input noise of the analyzer.
Where is the noise floor for both measurements?

Quote:

I measured the driver noise by making some equivalent resistors for each driver.

For the Emcore Laser / Newport 6000:

  • R_equiv = 5.55 Ohms @ 0.48 W
  • I soldered four 22 Ohm 1/4W resistors in the E shop in parallel to make this
  • R_measured = 5.7 Ohms

For the Covega Laser / ITC 510:

  • R_equiv = 3.75 Ohms @ 0.43 W
  • I soldered four 15 Ohm 1/4W resistors in the E shop in parallel to make this
  • R_measured = 3.9 Ohms

Here are some plots:

  1. Current Noise of each driver
  2. Frequency Noise (assuming the tuning coefficients in previous post) from this current noise

 

  51   Tue Jan 4 18:53:45 2011 DmassLaserLaserBeat Found

Here is the RIN of the Diodes, as measured on the HP35670A.

A second plot is included, which shows what we think is the Driver contribution to the RIN...

  1. I took each laser diodes P vs I curve to get dP/dI, then divided by P_laser to get dRin/dI
  2. This gave me 0.138 W/A and 0.26 W/A for the Emcore and Covega Laser Didoes, respectively
  3. This corresponds to a dRIN/dI of 3.5/Amp and 5.2/Amp (if what I described in 1 is OK)
  4. The contributions to RIN in the noisebudget from the driver noise are HIGHER than the actual RIN measured
  5. I think this is fine and it merely means that the transfer function of current noise into diode RIN is not the flat unity gain model implicitly assumed in 1 above
  6. Maybe a good measurement to do at some point is to T a (low noise) buffer onto the input current to the diode, and measure the coherence and transfer function between the driver current noise and the laser diode RIN
  7. Using each laser diode's P(I) curve to get the transfer function of current noise into RIN seems increasingly less bad at low frequencies, which is I would expect (no crazy low frequency features in the transfer function) - and it looks like we might be limited by the driver noise for low frequency RIN.
  8. If RIN and frequency noise are coherent, and we close a loop on frequency noise, this won't be a huge problem, since actuator noise is supressed by our (not yet existent) loop gain, and it should crush the current noise...the question, as ever, is "Can we obtain enough bandwidth to lower the PSD to what we define as an acceptable value"?

 

Happy New Year CryoLab.

 Updated (FIXED) second plot

  • I trust neither the purple high frequency curve, not the black noise floor curves. I think they are wrong from reflection and lab book reading. I believe the rest (for now)
Attachment 1: DiodeRin.pdf
DiodeRin.pdf
Attachment 2: DiodeRinNoiseBud.pdf
DiodeRinNoiseBud.pdf
  50   Tue Jan 4 18:12:48 2011 DmassLaserLaserDriver Noise Measured

I measured the driver noise by making some equivalent resistors for each driver.

For the Emcore Laser / Newport 6000:

  • R_equiv = 5.55 Ohms @ 0.48 W
  • I soldered four 22 Ohm 1/4W resistors in the E shop in parallel to make this
  • R_measured = 5.7 Ohms

For the Covega Laser / ITC 510:

  • R_equiv = 3.75 Ohms @ 0.43 W
  • I soldered four 15 Ohm 1/4W resistors in the E shop in parallel to make this
  • R_measured = 3.9 Ohms

Here are some plots:

  1. Current Noise of each driver
  2. Frequency Noise (assuming the tuning coefficients in previous post) from this current noise
Attachment 1: CurrentNoise.pdf
CurrentNoise.pdf
Attachment 2: FreqNoise.pdf
FreqNoise.pdf
  49   Tue Jan 4 17:12:47 2011 DmassLaserLaserLinewidth

Here is a picture of the beat measurement between the Emcore and Thorlabs laser diodes.

PD: Newport 818-BB-30 (2.5 GHz)

Analyzer: HP 8560E (~3 GHz?)

 

I used a couple points from the picture (via the analyzer cursor) to estimate the linewidth to be ~800 kHz (+/- 200 kHz). This is consistent with what Emcore told us the linewidth is.

  • I used 3 points and CRUDELY fit a lorentzian...this is why my error bars are huge.
  • I see what appears to be sidebands 2.3 MHz away from the center peak, 5.2 dB down (in amplitude)
  • Both lasers are (supposedly) single mode in their respective operating regions, which encompass the current and temperatures we were running them at according to their data sheets and an email to one of the Emcore engineers.
  • The sideband suppression ratio (SBSR) is quoted as > 30 dB for both of these lasers

At longer time scales, the laser frequency fluctuated a lot. Here are some videos:

Beat: 50 MHz/div and 10 dB/div (Youtube)

Beat: 10 MHz/div and 10 dB/div (Youtube)

The second video shows an uglier duckling.

 

 

 

Attachment 1: IMG_0376.JPG
IMG_0376.JPG
  48   Wed Dec 22 01:11:45 2010 DmassLaserLaserBeat Found

I found the beat between the two laser diodes!

What we knew about them before:

Covega (Thorlabs):

  • Operating Temp: 30 C (chip)
  • Operating Current: 340 mA
  • Thermistor is an NTC, the Thorlabs controller reads out resistance
    • R(30 C) = 8111 kOhms (see plot)
  • Lambda_measured ~ 1550.26 (with datasheet)

Emcore:

  • Operating Temp: 24.73 C
  • Operating Current: 301.1 mA
  • Lambda - 1550.12 (+/- unknown - has to be less than 0.4 nm by ITU specs, but probably better)

I made a roughly balanced, roughly mode matched setup on a fast PD (Newport 818-BB-30). (Pics of setup later).

  • I saw that tuning the Thorlabs LD via current might be insufficient if the diodes are more than 0.1 nm apart
  • I tuned the temperature to find the beat - it took some fiddling, but a couple Kelvin was sufficient.
  • I tried the following recipe:
    1. Put both Laser Diodes at their set Temp and Current
    2. Detune Laser Diode B's temperature to put the beat in the range of a scope/spectrum analyzer (2.5 GHz HP 8560E)
    3. Tune Laser Diode A's temperature / current to find its response
  • There was a problem - the Thorlabs (Covega) diode either mode hopped, or went into multi mode operation while detuning its temperature to find the beat with the Emcore LD
  • I chose to make the measurements with the Emcore LD 1 degree hot. This may increase linewidth, but it's something we can establish later
  • I used the scope for everything in it's band (and slightly out of)  to get better numbers for frequency

The measurement:

 

Laser Diode Beat
Emcore Emcore Covega (Thorlabs) Covega (Thorlabs)  
Temp (C) Current (mA) Temp (kOhms - 10k NTC) Current (mA) f_beat (MHz)
25.71 301.1 8.109 339.7 184
25.71 301.1 8.097 339.7 321
25.71 301.1 8.109 339.7 150
25.71 301.1 8.109 341.3 291
25.74 301.1 8.112 341.3 150
25.76 301.1 8.112 341.3 350
25.88 301.1 8.112 341.3 2000
25.74 301.1 8.112 341.3 150
25.74 301.4 8.112 341.3 340
25.74 302 8.112 341.3 700

This gives us:

Covega:

  • df/dT = 3.6 GHz / Kelvin
  • df/dI = 90 MHz / mA
  • (from datasheet) dlambda/dI  = 0.1 nm / 150 mA = 0.67 nm/A
  • The above => df/dI = dlamba/dI*(1.9*10^14 Hz/1550nm) = 86 MHz/mA => numbers in green check out.

Emcore:

  • df/dT = 13 GHz / Kelvin
  • df/dI = 630 MHz / mA

Next up:

  • "Linewidth"  of beat
  • Diode driver current noise
  • RIN of each PD
Attachment 1: CovegaNTC.pdf
CovegaNTC.pdf
Attachment 2: CovegaNTCRvsT.pdf
CovegaNTCRvsT.pdf
  47   Fri Dec 17 00:47:37 2010 DmassLaserLaserDiode Laser Measurements

I made the cables to go from the Newport 6000 to the LM14S2, (and had Frank check them).

We already checked the Thorlabs ITC510 and its cabling to the LM14S2 while measuring the JDSU Diode (Type 1).

To check the cabling I made to the Newport 6000, I put the JDSU diode into the LM14S2 (because it is the least bad one to explode).

  • The TEC worked fine (though there was some silly offset in there ~30 ohms)
    • setting "300 Fast", setpoint = 7.16kOhms
  • The laser diode seemed fine:
JDSU LD with Newport 6000
Current Power
40 mA 4.5 mW
80 mA 15.7 mW
110 mA 23.9 mW

NEXT - I swapped the Emcore LD with the JDSU (they have the same pinout), and mounted both laser diodes in collimators

Testing the Newport 6000 with the Emcore:

  • Set the TEC - same setpoint (10k NTC Thermistor again) - 7.16 kOhm - same settings - fine
  • Turned on the Emcore Laser (threshold 17 mA)
  • Current vs Power:
Emcore 1772 with Newport 6000
Current (mA) Power (mW)
15 0.006
17 0.010
19 0.174
21 0.236
23 0.614
30 1.82
50 5.20
75 9.30
100 13.3
150 21.3
200 28.9
250 36.0
260 37.6
270 39.12
275 39.9
280 40.5
285 41.2
290 41.8
292 42.0
294 42.25
296 42.5
298 42.7
300 42.90
301.1 43.02

 Next up was the Covega SFL-6881 (from Thorlabs): The Covega datasheet has a very nice I/V/Power plot (among others) - I will scan these at some point and post them. To confirm working-ness, Here are a couple points:

Covega (Thorlabs) SFL Diode
Current (mA) Power (mW)
   
245 40.1
275 47.9
  • Lasing threshold appeared about 50 mA (consistent with datasheet plot)
  • These numbers above seem high when compared with the Covega measurement.
  • The Covega measurement seems very professional, or at least their axes look nice.
  • Thermistor resistance set to 7.16k (around 30 C)
  • The Covega measurement was done at 30 C
  • When I fiddled with the temperature setpoint, the power change...but not so much...
  • I eyeballed the NTC Thermistor plot in the Newport 6000 Manual (page 74), and dR/dT ~(15kOhms/60 K) = 250 Ohms/C
  • ~100 Ohms (~0.4 Kelvin) change in the TEC setpoint caused a 1 mW change in power. There seems to be extra power on the order of 30% (48 mW where there "should" be ~37 mW).

Could our new Thorlabs power meter be off with a systematic of up to 30% in the 40 mW range? (This is the upper end of the range it's specced for)...

Set up: I put the Thorlabs power meter ~ 1inch in front of the collimator to capture ~ all the power.

 

  46   Wed Dec 15 18:30:05 2010 DmassLaserSchematicsDiode Laser Wiring

The diode lasers are all here:

We have two butterfly mounts from Thorlabs

The drivers we have / are borrowing are:

The butterfly mount (LM14S2) is the adapter between the butterfly packaged diode and the driver. There is some internal wiring in the LM14S2 which is diode / driver specific.

Laser Diode Chart - for reference
  Thorlabs Emcore JDSU

LM14S2 config

Type I Type 2 Telcom Laser Diode same as Emcore
1 TEC (+) Thermistor  
2 Thermistor Thermistor  
3 N/C Laser DC Bias (-)  
4 N/C MPD Anode  
5 Thermistor MPD Cathode  
6 N/C TEC (+, current in cools) TEC Anode
7 N/C TEC (-) TEC Cathode
8 N/C Case Ground  
9 N/C Case Ground  
10 Dev Anode No Connection  
11 Dev Cathode Laser Common (+) LD Anode (Case)
12 N/C Laser Modulation (-) LD Cathode (AC Input)
13 Case Laser Common (+) LD Anode (Case)
14 TEC (-) No Connection  

Diode Pinout:

7........1

DIODE  >==(tail)===

8.......14

The LM14S2 comes with a type 1/2 configuration card depending on which type of LD we are using it with.

For connecting the LM14S2 to the driver, the Thorlabs ITC510 had cables already made...

 

 

 

  45   Mon Dec 13 08:34:50 2010 FrankPhotosGeneralnew laser table

100_1165.JPG

100_1166.JPG

  44   Wed Dec 8 17:54:57 2010 FrankLab InfrastructureGeneralnew optical table will be moved tomorrow morning

i disassembled the existing stuff and moved the breadboard and it's base together with Jan and Tara within the lab out of the way to make space for the new table.
The laser diodes are back in it's original package.

The new table will be moved tomorrow morning 8am.

We marked the position of the table on the floor using electrical tape and already moved four legs to that room.

  43   Tue Dec 7 17:20:26 2010 FrankCryostatSensorsCryo sensor overview & stability vs thermal cycling

graphs taken from here: Sensor_Design_and_Operation_lakeshore.pdf

Sensor_Design_and_Operation.png

thermal_cycling_germanium.png

thermal_cycling_germanium2.png

  42   Tue Dec 7 17:10:32 2010 FrankCryostatSensorstypical cryo sensor controller stability vs temperature

 

taken from one of the Lakeshore temp controllers. Depending on the sensor 5-10mK are possible around 125K, even better at 18K

cryo_sensors_stability.png

  41   Sun Dec 5 19:52:58 2010 FrankLiquidGeneralEditing the ELOG

The link given is pointing to a wrong entry. First it is pointing to an entry to the Eager elog which does not exist anymore and second it does not describe how to edit the configuration.

Quote:

http://nodus.ligo.caltech.edu:8080/40m/3384 See this post in the 40m elog, it describes what files on nodus to edit to change things. Hack away!

DYM: The entry is fine, the information is simply out of date. The location of the files to edit (where are they on nodus) should be clear from the post. To figure out what to edit, either poke around (what I did), or rtm. I don't recall anything being not straight-forward, so I didn't record it for posterity.

  40   Tue Nov 30 20:18:45 2010 FrankHardwareSensorsGermanium temp sensors

This is what Lakeshore says: (http://www.lakeshore.com/temp/sen/gtrd.html)

"Germanium sensors have a useful temperature range of about two orders of magnitude. The exact range depends upon the doping of the germanium element.
Cryogenic temperature sensors with ranges from below 0.05 K to 100 K are available.
Between 100 K and 300 K, dR/dT changes sign and dR/dT above 100 K is very small for all models.
Sensor resistance varies from several ohms at its upper useful temperature to several tens of kilohms at its lower temperature."

  • High sensitivity provides submillikelvin control at 4.2 K and below
  • Excellent reproducibility better than ±0.5 mK at 4.2 K

So for the 18K zero CTE point germanium is probably the better choice.

 

Silicon diode for comparison:

Calibrated Accuracy
 

Typical sensor accuracy2

1.4 K
±12 mK
4.2 K
±12 mK
10 K
±12 mK
77 K
±22 mK
300 K
±32 mK
500 K
±50 mK
2 [(Calibration uncertainty)2 + (reproducibility)2] 0.5 for more information see Appendices B, D, and E
Long-term Stability
 

Use to 305 K3

Use to 500 K 4
4.2 K
±10 mK
±40 mK
77 K
±40 mK
±100 mK
305 K
±25 mK
±50 mK
500 K
±150 mK
3 Long-term stability data is obtained by subjecting sensor to 200 thermal shocks from 305 K to 77 K
4 Based on 670 h of baking at 500 K

 

  39   Mon Nov 29 16:41:10 2010 FrankGlassGenerallinear thermal expansion for Si and temp noise

i'm currently working on a reliable number for the thermal expansion of silicon to determine the temp stability required t
I pulled the data from the publications listed on the wiki. The first point of interest is around 124K.

The original data is from G.K. White from 1973 or so. But he does not provide a formula, only data points in a table which i was too lazy to copy number by number (it's a scanned copy, so no copy&paste :-() and then do the fit on my own. So i got the polynomial data from this: THE JPL CRYOGENIC DILATOMETER: MEASURING THE THERMAL EXPANSION COEFFICIENT OF AEROSPACE MATERIALS.
JPL measured their own samples and compared it to the previous measured ones to verify their measurements.

Here the function valid from 30K to 300K:

function varout = alpha_Si(T)
varout = -6.794490651331250E-7 ...
         +7.471762934256760E-8.*T ...
         -2.862997615244640E-9.*T.^2 ...
         +4.582058436059240E-11.*T.^3 ...
         -4.001740284246030E-13.*T.^4 ...
         +2.207261334864402E-15.*T.^5 ...
         -8.018923094512611E-18.*T.^6 ...
         +1.870326588771600E-20.*T.^7 ...
         -2.533566089676708E-23.*T.^8 ...
         +1.507992248919550E-26.*T.^9;

Here some plots:

alpha_Si.png

and zoomed in around 124K:

alpha_Si_125K.png

I assume that we can hold the temperature in a 100mK window around the zero CTE point at 124.33K. The good diode sensors have a  reproducibility ±10 mK at 4.2 K and a long-term stability of about 40mK. So i think 100mK is fair. The CTE at 124.28K is then -8.415e-10/K and at 124.38K it is 8.89e-10/K. So a good first number for a realistic expansion coefficient is something like 1e-9/K to be on the safe side.

Working on numbers for the lower zero CTE point.

On the other hand people did temp stabilization at cryogenic temperatures before and i found one description of the noise spectra (but no graph). The Lakeshore temp controllers (old model, DRC 91 CA) can do something like 30uHz/sqrt(Hz) flat with 1/f starting around 1e-2Hz.

  38   Mon Nov 29 16:20:11 2010 DmassHowToGeneralEditing the eog

http://nodus.ligo.caltech.edu:8080/40m/3384 See this post in the 40m elog, it describes what files on nodus to edit to change things. Hack away!

  37   Sat Nov 27 22:23:22 2010 FrankGlassPlotsWindow Transmission Curves & diagram from Oxford Instruments

Document from Oxford Instruments

Quote:

Window Transmissions from Janis

 

  36   Sat Nov 27 21:15:58 2010 FrankLiquidGeneralNFPA sign for WB room 050

NFPA_Sign_WB050.png

NFPA_Sign_WB050.pdf

  35   Wed Nov 24 16:47:12 2010 ranaGlassPlotsWindow Transmission Curves from Janis

Window Transmissions from Janis

  34   Tue Nov 23 18:58:29 2010 DmassGlassPlotsThe "window" problem

I didn't understand why windows might be a problem in terms of heat loading, so I made this plot!

Two questions:

  1. Does the (heavily attenuated) room temperature radiation from 400nm to 2um cause any problems?
    1. No - it gives a load of order 100 pW at hte cold stage assuming some reasonable geometries
  2. Does the emissivity of the windows matter?
    1. YES! If you assume some straight through beam path geometry, with windows on each radiation shield, I found approximately a 20% increase in thermal load from having 1" windows
Attachment 1: BBody.pdf
BBody.pdf
  33   Mon Nov 22 22:52:11 2010 DmassLaserGeneralWork has begun!

We started work on diode characterization today.

  • I mounted the JDSU CQF935 laser diode on the Thorlabs LM14S2 butterfly mount, again using Arctic Silver 5 CPU grease.
  • We looked at the manuals for the Thorlabs ITC510 temperature controller, and figured out how to set the dip switches on the back so that we could turn on the diode without blowing the world up.
  • We turned on the TEC and tuned it to the operating point given in the diode data sheet (7.16kOhms) - the sensor is a ~10k NTC
    • The readback of the loop was stable at the 1 Ohm level on the 7144 Ohm operating point, and looked stable on a scope (less than 1% rms)
  •  Removed the monitor PD bias on the ITC510 via the trim pot
  • Turned on the diode to measure P vs I (PLOT BELOW)
  • Checked that we had about 10 mW at 60 mA (we had 9.95)
  • Grabbed the Agilent 35670A to measure RIN spectra as a function of current (PLOT BELOW)
Attachment 1: JDSUPvsI.pdf
JDSUPvsI.pdf
Attachment 2: JDSUrin.pdf
JDSUrin.pdf
  32   Thu Nov 18 20:14:40 2010 FrankHardwareLaserButterfly arrived today

i picked up the second butterfly mount including three adapters for two different standard pinouts and one board for fully custom wiring. Whenever the first diode shows up we can get started. We should measure the diode driver noise until then to know what noise level we deal with. Once we measured the tuning coeff for the diodes we know the frequency noise contribution from the diode driver current fluctuations.

  31   Thu Nov 18 20:10:37 2010 FrankLaserPlotsDiode RIN

the end of the fiber is exactly how they sell it if you don't order it with a connector.They actually added kind of a strain relief to hold the end in some mount. The JDSU diode i've ordered will probably come with a real bare fiber end which is usually spliced to your equipment.

Quote:

Quote:

In an effort to get started, I am using a Lumics tunable diode laser from the 40m. I made some progress after the diode itself was eventually located on a bench in the PSL lab with no record of it having been signed out from the 40. I will rectify the lack of record tomorrow.

Initial cavity work at 1064 nm will take place in parallel with characterization of 1550nm diode lasers (as they get here.)

  • I had to remount the LUMICS PD in the Thorlabs butterfly package -
  • I used acetone then methanol to get the old grease off
  • I applied a thin layer of Arctic Silver 5 (good CPU paste) to the diode to stick it to the heatsink

The ferrule on the LUMICS diode is not entirely complete - it seems to be missing a threaded piece that would normally screw in to collimators etc. I measured laser RIN for different diode currents, and noticed that at higher currents, the corner where the RIN flattened out moved up in frequency.

I found an old elog from Sam which gives some previous measurements of this diode. Pictures to follow.

 Here are pictures of the diode:

  • before I cleaned its bottom with acetone / methanol
  • After I greased it up with Arctic Silver 5 cpu paste
  • Mounted in the Thorlabs heatsink

 

  30   Wed Nov 17 16:08:10 2010 DmassLaserPlotsDiode RIN

Quote:

In an effort to get started, I am using a Lumics tunable diode laser from the 40m. I made some progress after the diode itself was eventually located on a bench in the PSL lab with no record of it having been signed out from the 40. I will rectify the lack of record tomorrow.

Initial cavity work at 1064 nm will take place in parallel with characterization of 1550nm diode lasers (as they get here.)

  • I had to remount the LUMICS PD in the Thorlabs butterfly package -
  • I used acetone then methanol to get the old grease off
  • I applied a thin layer of Arctic Silver 5 (good CPU paste) to the diode to stick it to the heatsink

The ferrule on the LUMICS diode is not entirely complete - it seems to be missing a threaded piece that would normally screw in to collimators etc. I measured laser RIN for different diode currents, and noticed that at higher currents, the corner where the RIN flattened out moved up in frequency.

I found an old elog from Sam which gives some previous measurements of this diode. Pictures to follow.

 Here are pictures of the diode:

  • before I cleaned its bottom with acetone / methanol
  • After I greased it up with Arctic Silver 5 cpu paste
  • Mounted in the Thorlabs heatsink
Attachment 1: IMG_0334.jpg
IMG_0334.jpg
Attachment 2: IMG_0336.jpg
IMG_0336.jpg
Attachment 3: IMG_0340.jpg
IMG_0340.jpg
  29   Tue Nov 16 04:21:25 2010 DmassLaserPlotsDiode RIN

In an effort to get started, I am using a Lumics tunable diode laser from the 40m. I made some progress after the diode itself was eventually located on a bench in the PSL lab with no record of it having been signed out from the 40. I will rectify the lack of record tomorrow.

Initial cavity work at 1064 nm will take place in parallel with characterization of 1550nm diode lasers (as they get here.)

  • I had to remount the LUMICS PD in the Thorlabs butterfly package -
  • I used acetone then methanol to get the old grease off
  • I applied a thin layer of Arctic Silver 5 (good CPU paste) to the diode to stick it to the heatsink

The ferrule on the LUMICS diode is not entirely complete - it seems to be missing a threaded piece that would normally screw in to collimators etc. I measured laser RIN for different diode currents, and noticed that at higher currents, the corner where the RIN flattened out moved up in frequency.

I found an old elog from Sam which gives some previous measurements of this diode. Pictures to follow.

  28   Mon Nov 15 15:28:41 2010 FrankShoppingGeneral20mW PM fiber coupled laser diode CQF935/508 ordered

I've ordered the 20mW PM fiber coupled laser diode CQF935/508 with my credit card. Should be here within the next days...

  27   Mon Nov 15 15:27:05 2010 FrankShoppingPlotsCavity Higher Order Mode Spacing

is yellow for 1m curvature? the legend for the fused silica cavity says 50cm for both...

Quote:

Here is a plot of the HOM spacing for the Al and fused silica cavities for curved / curved, with R=50cm and R=100 cm

 

The height of the modes (colored x's) corresponds to their order (N+M) for TEM_{NM}

 

 

  26   Mon Nov 15 13:31:43 2010 DmassShoppingPlotsCavity Higher Order Mode Spacing

Here is a plot of the HOM spacing for the Al and fused silica cavities for curved / curved, with R=50cm and R=100 cm

 

The height of the modes (colored x's) corresponds to their order (N+M) for TEM_{NM}

 

Attachment 1: CavityG-factors.pdf
CavityG-factors.pdf
  25   Wed Nov 10 14:08:47 2010 DmassShoppingGeneralDiode Laser Delay Update

I spent some time on the phone with Thorlabs and Emcore about some unrelated shenanigans happening on both orders.

For the Thorlabs diode, apparently our (CiT) policy is to just pay within 10% of the purchase order with not many questions asked. This differs from Thorlabs policy which is to request a new PO if there is a 100$ difference. The PO was stuck between CiT and Thorlabs with us saying "take the darn PO, we'll pay it, we can't waste our time for your book keeping", and Thorlabs saying "We can't do that, send us a new one".

I poked and prodded us and them, and Thorlabs promised to resolve it tomorrow morning (when someone's manager got back) and overnight it.

MORAL: Always overestimate the costs via techmart by rounding up to be safe.

 

Emcore - there was a 7-8 Week lead on what we requested (1772-C01-40-34...), so I talked to the engineer Frank was in contact with, and some lady in sales, and Dec 3rd is the soonest they say that they could get something shipped from their overseas factory - which would mean ~Dec 10th for us.

They were willing to start the ball rolling on getting us a 1772-40-34- (1550 nm 40 mW) made before they actually have a PO from us. We need to tell them in the next couple days whether we still want this. It will be 1k.

 

JDSU: In light of the Emcore delays, I think we should consider getting http://www.acronymeo.com/cqf935508.html this 20 mW used JDSU low linewidth diode. We would need to connectorize it, but Frank seems to know how to do that, and I think I would like to see how it's done. It is 500$.

  24   Wed Nov 10 02:57:35 2010 DmassLaser To dizzle

MATLAB CODE for cavities - cut and pastable

function [Preflnorm,Pcouplnorm,Ptransnorm,Fin] = FPcavity(varargin)
% This code gives the ratios of the electric fields in a resonant Fabry
% Perot etalon as a function of individual cavity mirror reflectivity and
% loss:
%
%     varargin = [Refl1,Refl2,L1,L2]
% Where R is the reflectivity (in power) of the mirrors (1 is front), and
% L is the power loss in ppm for each mirror
%  davidym@caltech.edu   -  11/9/2010

%% Define the variables to use for the fields
r1=sqrt(varargin{1});
r2=sqrt(varargin{2});
L1=varargin{3};
L2=varargin{4};
grt=r1.*r2.*sqrt((1-L1/1e6).*(1-L2/1e6));
t1=sqrt(1-r1.^2-L1/1e6);
t2=sqrt(1-r2.^2-L2/1e6);

Ereflnorm=1./r1.*(r1.^2-grt)./(1-grt);
Preflnorm=Ereflnorm.^2;
Ecouplnorm=t1./(1-grt);
Pcouplnorm=Ecouplnorm.^2;
Etransnorm=t1.*t2./(1-grt);
Ptransnorm=Etransnorm.^2;
Fin=pi*sqrt(grt)./(1-grt);

  23   Mon Nov 1 00:55:22 2010 ranaLaserGeneralDiode Laser Freq Noise Requirement

To be quantitative about the laser noise, lets adopt 2 possible servo configurations:

1) UGF = 500 kHz. gain increases like 1/f down to 50 kHz. Below 50 kHz increases like 1/f^2 all the way to DC (i.e. we have two true integrator stages).

2) UGF = 1 MHz. gain increase as 1/f down to 100 kHz. Goes up like 1/f^3 below 50 kHz.

 

#1 is much like the usual FSS servos in LIGO.

#2 is similar to the LIGO MC servo, but with ~10x higher bandwidth.

 

We should estimate the in-loop noise using both models to determine how aggressive we have to be in the servo design.

  22   Sun Oct 31 23:15:59 2010 DmassLaser Diode Laser Freq Noise Requirement

I added some traces to the plot so that it wasn't quite so useless.

Diodes 1 2 and 3 on the plot are pulled from the linked wiki page and translated into Hz/rtHz. I believe the linewidths on these lasers were ~1 MHz.

I do not know a bunch of high frequency excess laser noise would mean for us...to quantify:

If we have a lot of frequency noise at high frequency above the cavity pole, and the contributions to the rms frequency noise are dominated by the high frequency noise, I am unsure what that does to our error point and cavity transmission.

Attachment 1: NoiseBudDiodes.pdf
NoiseBudDiodes.pdf
  21   Tue Oct 26 00:09:13 2010 DmassShoppingPaper TrailLaser Diodes

A few more things to techmart -

1550 laser sign

Kentek goggles

  20   Sat Oct 23 17:31:00 2010 DmassLaserR+DDiode Laser Freq Noise Requirement

I recreated the elusive plot of diode laser frequency noise requirements which Frank has been referring to with the following recipe:

  1. I took the coating thermal noise limit on page 6 of the Cryo Cavity proposal (figure 3)
  2. Divided this spectrum by 10
  3. Assumed a PDH servo for locking to the cavity with a UGF of 500kHz, and a boost starting at 50 kHz
  4. Divided CTN/10 by the above closed loop transfer function
  5. Plotted it (Note I extended past 10^4 Hz)

For comparison, see the plot on the wiki under "diode lasers" - it's in dBMHz^2/Hz, which Rana and Frank think is a amplitude dB even though it's a PSD.

If its amplitude dB

  • -20 dB = 0.1 dBMHz^2/Hz => 1/3 MHz/rtHz
  • -40 dB = 1e-2 dBMHz^2/Hz => 100 kHz/rtHz
  • -60 dB = 1e-3 dBMHz^2/Hz => 33 kHz/rtHz

If its power dB (I'm not 100% convinced this is not the case), we have:

  • -20 dB = 1e-2 dBMHz^2/Hz => 100 kHz/rtHz
  • -40 dB = 1e-4 dBMHz^2/Hz => 10 kHz/rtHz
  • -60 dB = 1e-6 dBMHz^2/Hz => 1 kHz/rtHz

 The lowest frequency noise trace on the linked plots is the 60 mA CQF935 - it is -40 dBMHz^2/Hz at around 30 Hz. This is either 10 kHz or 100 kHz /rtHz, which is either "O.K" for our loop, or just too high to be suppressed by the (crude) loop we have assumed, at 30 Hz. THIS IS NOT NECESSARILY REPRESENTATIVE OF THE LASERS WE WILL BUY.

To do:

  • Infer a characteristic noise shape to the DFB laser from the linked curves
  • Calculate linewidths from modeled PSDs
  • Figure out what to do (how to treat it) if the high frequency noise really is white
  • Worry about what the high rms contributions from the unsuppressed high frequency laser noise do at our error point / to our transmitted beam

 

 

 

Attachment 1: DiodeFreqNoise.pdf
DiodeFreqNoise.pdf
ELOG V3.1.3-